OpenStudy (anonymous):

Perform the indicated operation. (-9 + 2i) - (-12 + 4i) =

OpenStudy (mathstudent55):

You can drop the first set of parentheses because they are unnecessary. To get rid of the second set of parentheses, you need to distribute the negative sign. Example: -(2x - 6) = -2x + 6

OpenStudy (mathstudent55):

Then add like terms.

OpenStudy (anonymous):

so its 3-2i?

OpenStudy (anonymous):

can i keep you for a few more questions?

OpenStudy (mathstudent55):

Correct.

OpenStudy (mathstudent55):

(-9 + 2i) - (-12 + 4i) = = -9 + 2i + 12 - 4i = -9 + 12 + 2i - 4i = 3 - 2i

OpenStudy (anonymous):

The product of a complex number and its conjugate is always A) real. B) imaginary. C) rational. D) natural.

OpenStudy (mathstudent55):

Do you know what the conjugate of a complex number is?

OpenStudy (anonymous):

nope

OpenStudy (mathstudent55):

Ok, let's start there.

OpenStudy (mathstudent55):

For a complex number a + bi, its complex conjugate is a - bi. This is very simple. The complex conjugate of a complex number is another complex number in which all you do is change the sign of the imaginary part (the part with the i).

OpenStudy (anonymous):

so the answer i its real

OpenStudy (anonymous):

is

OpenStudy (mathstudent55):

Examples: Complex Number Complex Conjugate 5 + 3i 5 - 3i -4 + 2i -4 - 2i 6 - i 6 + i -2 - 3i -2 + 3i 5 5

OpenStudy (mathstudent55):

Correct, but do you know why?

OpenStudy (anonymous):

because adding to real numbers together always give you real numbers?

OpenStudy (mathstudent55):

You're not adding them. You are multiplying a complex number and its conjugate.

OpenStudy (anonymous):

ok cool

OpenStudy (anonymous):

Select the BEST classification for π. A) irrational Eliminate B) rational C) imaginary D) complex

OpenStudy (anonymous):

is this one irrational

OpenStudy (mathstudent55):

Let's use the general complex number a + bi. Let's even say that b is not zero. this way the number does have its imaginary part. The complex conjugate of a + bi is a - bi. Now we multiply a + bi by its complex conjugate: (a + bi)(a - bi) Notice that this is the product of a sum and a difference. It follows the pattern: \((x + y)(x - y) = x^2 - y^2\) (If you forgot this pattern, you can always use FOIL and collect like terms.) We can use the pattern. \(a + bi)(a - bi)\) \( = a^2 - (bi)^2 \) \(= a^2 - b^2i^2\) \(= a^2 - i^2b^2\), but \(i^2 = -1\), so \(= a^2 - (-1)b^2\) \(= a^2 + b^2\) As you can see there is no i in the final expression, so \(a^2 + b^2\), the product of a complex number and its complex conjugate, is a real number.

OpenStudy (mathstudent55):

Correct, \(\pi\) is irrational.

OpenStudy (anonymous):

oh it makes sence now

OpenStudy (anonymous):

(x 1 a )(x 1 b ) Write the expression in simplified radical form.

OpenStudy (mathstudent55):

Great.

OpenStudy (mathstudent55):

Sorry, gtg.

OpenStudy (anonymous):

oh ok thank you!!